# Bad Students Make Great Teachers: Active Learning Accelerates Large-Scale Visual Understanding

### v2

@@ 4 Experiments
All our experiments were conducted with Vision Transformers [8] for which strong baselines are available across model sizes [44]. Unless specified, we adopt models with patch size 16 throughout (ViT-S refers to ViT-S/16 and similar). Since we are interested in studying active learning for large-scale pre-training, we consider two canonical regimes: large-scale classification on JFT-300M [37] and multimodal contrastive learning [30]. When pre-training with JFT classification we use held-out classification performance as the evaluation metric. When pre-training on image-text data we evaluate with standard multimodal transfer tasks: ImageNet zero-shot classification and image-to-text / text-to-image retrieval on COCO.
Throughout, we will refer to the large batch of size $B$ sampled IID from the training data as the ‘super-batch’, and the prioritised smaller batch of size $b<B$ as the ‘sub-batch’. In all our experiment, we filter 50% of IID sampled data such that $\rho=B/b=2$, although more aggressive filtering regimes warrant investigation [36].
Our models were developed using the JAX ecosystem [5, 4] with efficient prioritized replay based on [6].
@@ 4.1 Evaluating loss-based scoring heuristics in the large-data regime
We begin by evaluating a suite of loss-based scoring heuristics on their ability to accelerate supervised classification on the JFT dataset. Figure 3 shows the results of training an 86M parameter ViT-B with various sampling strategies. Arguably the most intuitive method to score data is to prioritise training on data with high loss under the learner. This has been shown to be effective in the small data regime [23, 11] but has not been successfully scaled. In our experiment, this hard learner prioritisation (equation 2) marginally improved performance at the end of training but was not significantly better than the IID baseline, despite requiring an additional inference pass over the super-batch. This is perhaps not surprising - data points with high loss may also be unlearnable due to e.g. label noise, such that training on those data points does not result in the model performing any better on the held out test set. Large scale datasets are more likely to be noisy.
Figure 3: Evaluation of loss-based data-selection criteria for large-scale classification. We train a ViT-B on JFT-300M with different data-selection policies. Prioritising hard data under the learner (green curve) produced late and marginal gains over the IID baseline. Prioritizing data using both learnability (blue curve, [28]) and ease with respect to a pre-trained reference model (red curve, [13]) produced significant speedups and performance gains.
@@
Scoring methods based on pre-trained reference models performed much better—both easy reference prioritisation (equation 3), and prioritising based on learnability (equation 4) produced significant gains over IID sampling. Here, we pre-trained an identical ViT-B for the same 3 epochs to use as a reference model for a second training run displayed above. In our experiments, learnability based scoring slightly underperforms easy reference despite requiring an extra inference pass on the learner. Both models produced speed-ups relative to IID training of  33% (Fig. 3).
4.2 Generalising data-selection policies across scale
Figure 4: Generalization of data-selection policies across models scales. Left: We train a ViT-L for 3 epochs on JFT using IID sampling (grey) or prioritized data sampling using example learnability (blue) or example difficulty (red). Example scores are computed using ViT-B actors (dark), or cheaper ViT-S or ViT-Tiny models (light). While both example learnability and difficulty yield good speedups with expensive actors, learnability criteria are much more robust to approximate scoring. Top right: Learner speedup is computed as the fraction of learner iterations saved in order to attain the baseline’s top performance. Actor overhead is computed as the the additional computation in FLOPs required to score examples with a particular actor architecture. Example learnability yields robust learner speedups across actor scales, example difficulty does not. Lower right: total compute efficiency is calculated as a product of learner efficiency and actor overhead, indicating the amount of computation required to reach the baseline’s performance. Approximate actors (i.e. ViT-S or smaller) computing example learnability enable total compute speedups, other schemes do not.
Although the speed-ups afforded by learnability prioritization (Fig. 3) are significant in terms of learner FLOPs, they come at the cost of the additional inference passes required to score the data during learner training, plus the cost of training the reference model. This makes the overall gains strongly compute-negative relative to IID training (in addition to requiring extra time for reference model pre-training). Even if the size of the reference model is scaled down [28], these methods still incur the cost of additional learner inference passes to score the data during training.
As discussed (Section 3.3), although easy-reference scoring could be amortized by constructing ‘foundation datasets’, the same savings could not be recovered in the semi-infinite data regime where data may never be repeated, or when training continually larger models.
Unlocking compute-positive active learning. To address this issue, we introduce a set of down-scaled models with the same ViT architecture that we use to score data for training a larger ViT-L model (see Methods Section 3.3). Specifically, we use ViT-B, ViT-S or ViT-Ti variants (which are  4$\times$, 13$\times$ and 47$\times$ cheaper than the learner model) for both the online and reference model (see Algorithm 1). In Figure 4 (left) we assess the impact of these cheaper scoring models on learner efficiency. First, we find that loss-based prioritization (i.e. easy reference) to be very sensitive to the capacity of the scoring model: while ViT-B scoring models yield reasonable gains over IID sampling, prioritizing with ViT-S and ViT-Ti scoring models underperforms significantly (Fig. 4, red curves).
In contrast, we find prioritized sampling according to example learnability to yield robust gains, even when the scoring models are significantly scaled down (Fig. 4, blue curves). For example, while ViT-B scoring models yield a 31% learner speedup, the 50$\times$ smaller ViT-Ti scoring models still provide a 26% speedup. We pushed this logic by using even smaller scoring models (the ViT-Mu family which we introduce, see Appendix A.7) which are up to 1000$\times$ smaller than the learner. Despite this, prioritizing data based on their scores yields non-negligeable speedups (e.g. 16% for the smallest actors we consider; Figure 4, top right).
These experiments demonstrate that, with the appropriate scoring criterion, online and reference models can be significantly downscaled and still produce comparative gains to the larger models, with learner efficiency degrading gracefully with the actor overhead (i.e. the cost of the reference model and data scoring). As a result, our method ClassAct quickly becomes FLOP positive as the online + reference models are downscaled (Figure 4, bottom right), while at the same time producing speed-ups in wall-clock time for a given learner batch size.
@@ 4.2 Generalising data-selection policies across scale
Figure 5: Scaling laws for active learning. To test the extent that smaller models can be used to accelerate pre-training for larger models, we trained a baseline ViT-L over a range of compute budgets (this is the compute optimal model size, see Zhai et al., 2021). We also trained the same ViT-L with both ViT-Ti and ViT-S reference policies, pre-trained for the same number of epochs. Left: Small model policies produce robust savings in learner compute. Right: When accounting for total compute (learner + actor training and data scoring), small model policies in all compute budgets produce FLOP savings over IID training. These scaling laws generalize those measured empirically in the IID setting [44] to the case of non-IID data selection.
4.3 Generalising neural scaling laws to the active-learning setting
We next investigated the scaling behaviour of ClassAct by experimenting over large learner compute budgets (Fig. 5), using both ViT-Ti and ViT-S models as actors for training a ViT-L. Predictably, the ViT-S produced larger, although marginal gains over the ViT-Ti actors when not accounting for scoring FLOPs (Figure 5, left). However, when accounting for total FLOPs, the difference was less pronounced (Figure 5, right). Our results generalize large scale IID scaling laws such as uncovered by [19] for LLMs and reproduced for large vision transformers by [44] to the case of non IID sampling. For the first time, we demonstrate that these scaling laws can be shifted in our favour by selecting data using general model-based scoring heuristics.
@@ 4.3 Generalising neural scaling laws to the active-learning setting
ViT-B IID B 0% 0%
RHO [28] Tiny B B Held-out, fixed 0% - 79%
@@
Table 1: Simplifying and accelerating the computation of learnability scores. Relative to RHO ClassAct  makes two changes: replacing the reference model with one trained in-domain (removing the need for bespoke held-out sets), and dramatically reducing capacity of the online actor models used for scoring examples. All experiments were conducted on 3 epochs of one-half of the JFT dataset to enable the held-out ablations. RHO with a small reference model did not produce a learner speedup in our experiments.
@@ 4.4 Training the reference model in parallel
We confirmed our hypothesis by running an experiment in which we trained our reference model on a super-batch of size $10b$ and trained the online and learner model in sequence with the sub-batch of size $b$. To make sure the reference model quickly converges, we additionally set the learning rate to double that of the online and learner models (this would cause instability for the learner and online models because of the additional variance from the smaller batch). We also verified that training the reference model on a held-out set of data performed equally in our experiments to reference models trained on the same data as the learner model [28]. Our ‘one-pass’ setup, Online-ClassAct, produces the same performance as the pre-trained ClassAct pipeline in our Ti-trains-B experiments (Table 1). Pseudocode is shown in appendix Algorithm 2.
@@
ImageNet 0-Shot COCO (im2txt) COCO (txt2im)
Reference Learner (+online) Speed-up Top 1 Accuracy Speed-up R@1 Speed-up R@1
IID ALIGN 0.0% 45.6 0.0% 38.5 0.0% 25.5
ActiveCLIP ALIGN ALIGN 23% 47.8 28.6% 40.9 26.8% 27.3
ActiveCLIP LTIP ALIGN 48% 53.2 48.0% 44.8 27.0% 30.5
IID LTIP 0.0% 46.6 0.0% 44.4 0.0% 30.1
ActiveCLIP ALIGN LTIP 0.0 % 46.5 18.8% 45.6 28.2% 31.8
ActiveCLIP LTIP LTIP 16% 47.2 22.1% 45.4 17.5% 31.0
Table 2: Generalizing data-selection policies across datasets and tasks. We pretrain reference models on the large, noisy ALIGN dataset, or the smaller and more curated LTIP dataset [2]. Consistently with [2], we find IID training on LTIP to yield stronger transfer learning results than pre-training on ALIGN. Interestingly, these reference models can be used very effectively for data-selection on both LTIP and ALIGN, whereas ALIGN-pretrained reference models yield more modest speedups. All models are provided with 800M training images at resolution 128$\times$128, speedups are shown relative to the time at which the IID baseline was reached for that evaluation metric.
@@ 4.5 ActiveCLIP: active multimodal learning
Figure 6: Reference policies generalize across tasks. We train a ViT-B reference policy on either ALIGN or LTIP ([2]), then use it to train a second ViT-B learner again on either ALIGN or LTIP. As before, ActiveCLIP produces efficiency gains over IID in terms of learner updates. The biggest gains, which surpassed the in-domain ActiveCLIP results, were instead found in the out-of-domain setting where a LTIP reference model is used to accelerate an ALIGN training run. See Table 2 for more details.
@@
Table 3: Comparison of ActiveCLIP to public multimodal pre-taining methods. All models use ViT-B/16 encoders, with 224$\times$224 image resolution, except SigLIP which uses 256$\times$256. ActiveCLIP outperforms models trained with the same or more training data, on ImageNet zero-shot classification and both forms of COCO retrieval. ${}^{\dagger}$benefits from additional ImageNet21K pretraining (with 2B training examples).
4.6 Policy generalization across tasks
To fully leverage off-the-shelf pre-trained image models for training (Figure 2), our results up to now suggest that the reference model should be trained on the same task as the learner model is being trained for. In Figure 6 (see Table 2 for full results), we show that we can in fact pre-train our reference models on related but distinct datasets. Moreover results suggest that there may even be a benefit to cross-training in some-cases. Models trained on ALIGN with an ALIGN trained reference model (‘in-domain’ Active-CLIP) produced similar speedups over models trained IID on ALIGN on downstream metrics, but these gains were greatly surpassed by instead training on the LTIP dataset. One possibility is that LTIP is less noisily labeled such that the scoring policies it produces are ‘cleaner’ - i.e. more able to filter for ‘clean’ data. The corollary may also be true; active data selection appears to greatly improve the utility of ALIGN, suggesting that it contains a large proportion of data that is of low quality for training.
4.7 Comparison to prior multimodal art
We leverage the insight that pre-training a reference model on clean data can facilitate learning on larger, noisy data for training our final model. Here, we train a reference model on LTIP, then use it to train a new model on the much larger mixture of LTIP and ALIGN, following [2], for a total of 3 and 8 billion training examples. Table 3 shows that in this training regime, ActiveCLIP surpasses models trained with significantly more data on ImageNet 0-Shot classification and COCO retrieval.
@@ 5 Discussion
In this work, we have presented a new method for active data selection that builds upon and simplifies the concept of ‘learnability’. Our experiments demonstrate that this approach can significantly reduce the computation required for large-scale pretraining, compared to uniform IID training. To our knowledge, this is the first active learning method that is more efficient than IID training when accounting for total FLOPs, and that does not rely on hand-designed features, allowing broad application across training setups. We have validated this by showing results on classification and contrastive pre-training, and found that our data selection policies continue to produce efficiency gains in the large-scale regime and can generalize effectively across task modalities. Collectively, our experiments also illustrate a Pareto frontier that allows trading off actor/data-selection computation against savings in training iterations, suggesting an alternative path to improved performance beyond scaling training batch sizes.
In this work we focused on supervised pre-training for images, but further work could involve extending our method to other modalities and training schemes such as language, video, and generative modeling. An important note is that all our experiments present results from filtering only 50% of the data; further gains may be possible by filtering more aggressively, at the cost of further overheads. In particular, aggressive data-selection coupled with efficient scoring schemes such as the ones proposed here could test the hypothesis that large-scale pretraining can benefit from exponential, rather than power-law, scaling behavior.
Appendix A Appendix: supplementary methods
A.1 Details of total compute calculations

### v4

@@ 4 Experiments
All our experiments were conducted with Vision Transformers [9] for which strong baselines are available across model sizes [43]. Unless specified, we adopt models with patch-size 16 throughout (ViT-S refers to ViT-S/16 and similar). We consider two canonical tasks for large-scale pretraining: classification on JFT-300M [36] and multimodal contrastive learning [30] on large image-text datasets. When pre-training with JFT classification we use held-out classification performance as the evaluation metric. When pre-training on image-text data we evaluate with standard multimodal transfer tasks: ImageNet zero-shot classification and image-to-text / text-to-image retrieval on COCO.
Figure 3: Evaluation of loss-based data-selection criteria for large-scale classification. We train a ViT-B on JFT-300M with different data-selection policies. Prioritising hard data under the learner (green curve) produced marginal gains over the uniform sampling baseline. Prioritizing data using both learnability (blue curve, [27]) and easy reference prioritization (red curve, [15]) produced significant speedups and performance gains.
Figure 4: Generalization of data-selection policies across models scales. Left: We train a ViT-L for 3 epochs on JFT using uniform sampling (grey) or prioritized data sampling using example learnability (blue) or low-loss under the reference model (red). Example scores are computed using ViT-B actors (dark), or cheaper ViT-S or ViT-Tiny models (light). While both example learnability and “easy reference” yield good speedups with expensive actors, learnability criteria are much more robust to approximate scoring. Top right: Learner (ViT-L) speedup is computed as the fraction of learner iterations saved in order to attain the baseline’s top performance. Actor overhead is computed as the additional computation in FLOPs required to score examples with a particular actor architecture (varying from ViT-Mu to ViT-L, see Appendix Table 4). Example learnability yields robust learner speedups across actor scales, “easy reference” scoring does not. Lower right: total compute efficiency is calculated as a product of learner efficiency and actor overhead, indicating the amount of computation required to reach baseline performance. Approximate actors (i.e. ViT-S or smaller) computing example learnability enable total compute speedups, other schemes do not.
Throughout, we will refer to the large batch of size $B$ sampled uniformly from the training data as the ‘super-batch’, and the prioritised smaller batch of size $b<B$ as the ‘sub-batch’. In all our experiment, we filter 50% of uniformly sampled data such that $\rho=B/b=2$, although more aggressive filtering regimes warrant investigation [35].
@@ 4.1 Evaluating loss-based scoring heuristics in the large-data regime
We begin by evaluating loss- and learnability-based heuristics on their ability to accelerate supervised classification on JFT (Fig. 3). Arguably the most intuitive method to score data is to prioritise training on data with high loss under the learner (hard learner). In our experiments, this strategy (Eq. 1) only marginally improved performance over the uniform sampling baseline, despite requiring an additional inference pass over the super-batch. This is perhaps not surprising - data points with high loss may also be unlearnable due to e.g. label noise, such that training on those data points does not result in the model performing any better on the held-out test set. Large scale datasets are more likely to be noisy.
@@
Scoring methods based on pre-trained reference models performed much better—both easy reference (equation 2) and learnability (equation 3) -based prioritization produced significant gains over uniform sampling. Here, we pre-trained an identical ViT-B for the same 3 epochs to use as a reference model for a second training run displayed above. producing speed-ups of  33% (Fig. 3).
@@
Figure 5: Scaling laws for active learning. We trained a baseline ViT-L over a range of compute budgets (for which ViT-L is compute optimal, see Zhai et al., 2021). We also trained the same ViT-L with both ViT-Ti and ViT-S reference policies, pre-trained for the same number of epochs. Left: Small model policies produce robust savings in learner compute. Right: When accounting for total compute (learner + actor training and data scoring), small model policies in all compute budgets produce FLOP savings over training with uniform samples. These scaling laws generalize those measured empirically in the uniform sampling setting [43] to the case of non-uniform data selection.
@@
Uniform Sampling (ViT-B) B 0% 0%
RHO Tiny B B Held-out, fixed 0% - 79%
@@
Table 1: Simplifying and accelerating the computation of learnability scores. Relative to RHO [27], ClassAct  makes two changes: replacing the reference model with one trained in-domain (removing the need for bespoke held-out sets), and dramatically reducing capacity of the online actor models used for scoring examples. All experiments were conducted on 3 epochs of one-half of JFT to enable the held-out ablations. RHO with a small reference model did not produce a learner speedup in our experiments.
4.2 Generalising data-selection policies across scale
The speed-ups afforded by learnability prioritization (Fig. 3) come at the cost of the additional inference passes required to score the data during learner training, plus the cost of training the reference model. This makes the overall gains strongly compute-negative relative to training with uniform samples. Even if the size of the reference model is scaled down [27], these methods still incur the cost of additional learner inference passes to score the data during training.
Unlocking compute-positive active learning. To address this issue, we introduce a set of down-scaled models with the same ViT architecture that we use to score data for training a larger ViT-L model (see Methods Section 3.3. We use ViT-B, ViT-S or ViT-Ti variants (which are  4$\times$, 13$\times$ and 47$\times$ cheaper than the learner) for both the online and reference model (see Appendix Algorithm 1). In Figure 4 (left) we assess the impact of these cheaper scoring models on learner efficiency. First, we find that easy reference prioritization to be very sensitive to the capacity of the scoring model: while ViT-B scoring models yield reasonable gains over uniform sampling, prioritizing with ViT-S and ViT-Ti scoring models underperforms significantly (Fig. 4, red curves).
In contrast, we find that learnability based prioritization yields robust gains, even when the scoring models are significantly scaled down (ClassAct; Fig. 4, blue curves). For example, while ViT-B scoring models yield a 31% learner speedup, the 50$\times$ smaller ViT-Ti scoring models still provide a 26% speedup. We pushed this logic by using even smaller scoring models (the ViT-Mu family which we introduce, see Appendix ) which are up to 1000$\times$ smaller than the learner. Despite this, prioritizing data based on their scores yields non-negligible speedups (e.g. 16% for the smallest actors we consider; Figure 4, top right).
These experiments demonstrate that, with the appropriate scoring criterion, online and reference models can be significantly downscaled and still produce comparable gains to larger models, with learner efficiency degrading gracefully with the actor overhead (i.e. the cost of the reference model and data scoring). As a result, our method ClassAct quickly becomes FLOP positive as the online + reference models are downscaled (Figure 4, bottom right), while at the same time producing speed-ups in wall-clock time for a given learner batch size.
4.3 Generalising neural scaling laws to the active-learning setting
We next investigated the scaling behaviour of ClassAct by experimenting over large learner compute budgets (Fig. 5), using both ViT-Ti and ViT-S models as actors for training a ViT-L. Predictably, the ViT-S produced larger, although marginal gains over the ViT-Ti actors when not accounting for scoring FLOPs (Figure 5, left). However, when accounting for total FLOPs, the difference was less pronounced (Figure 5, right). Our results generalize large scale uniform sampling scaling laws such as uncovered by [19] for LLMs and reproduced for large vision transformers by [43] to the case of non-uniform sampling. For the first time, we demonstrate that these scaling laws can be shifted in our favour by selecting data using general model-based scoring heuristics.
@@ 4.4 Training the reference model in parallel
We confirmed our hypothesis by running an experiment in which we trained our reference model on a super-batch of size $B=10b$ and trained the online and learner model in sequence with the sub-batch of size $b$. To make sure the reference model quickly converges, we additionally set the learning rate to double that of the online and learner models (this would cause instability for the learner and online models because of the additional variance from the smaller batch). We also verified that training the reference model on a held-out set of data performed equally in our experiments to reference models trained on the same data as the learner model [27]. Our ‘one-pass’ setup, Online-ClassAct, produces the same performance as the pre-trained ClassAct pipeline in our Ti-trains-B experiments (Table 1). Pseudocode is shown in appendix Algorithm 1.
@@
ImageNet 0-Shot COCO (im2txtR1) COCO (txt2imR1) (Learner)
Speed-up % (vs. Uniform) 23.0 48.1 28.6 48.0 26.8 27.0 ALIGN
0.0 16.3 18.8 22.1 28.2 17.5 LTIP
Performance (vs. Uniform) 47.8 (2.2) 53.2 (7.4) 40.9 (2.4) 44.8 (6.3) 27.3 (1.8) 30.5 (5.0) ALIGN
46.5 (-0.1) 47.2 (0.6) 45.6 (1.2) 45.4 (1.0) 31.8 (1.7) 31.0 (0.9) LTIP
(Reference) ALIGN LTIP ALIGN LTIP ALIGN LTIP
Table 2: Generalizing data-selection policies across datasets and tasks. We pretrain reference models on the large but noisy ALIGN dataset, or the smaller and more curated LTIP dataset [2]. Consistent with [2], we find training with uniform sampling on LTIP to yield stronger transfer learning results than pre-training on ALIGN. These reference models can be used very effectively for data-selection on both LTIP and ALIGN, whereas ALIGN-pretrained reference models yield more modest speedups. All models are provided with 800M training images at resolution 128$\times$128, speedups are shown relative to the time at which the uniform sampling baseline was reached for that evaluation metric. Colour indicates performance relative to uniform (brackets).
Figure 6: Reference policies generalize across tasks. We train a ViT-B reference model on either ALIGN or LTIP, then use it to train a second ViT-B learner on ALIGN with ActiveCLIP. The biggest gains were found with an LTIP reference model, despite it needing to perform out-of-domain generalization. In 50k iterations, ActiveCLIP selects the 800M “cleanest” examples from the ALIGN dataset, whose size is 1.6B in total.
@@ 4.5 ActiveCLIP: active multimodal learning
Figure 7: Reference models trained on curated datasets are powerful data selectors. Active data selection using ActiveSigLIP with reference models trained on increasingly curated subsets of Webli. Webli-350M reference models are consistently more effective than those trained on raw or 1B subsets.
4.6 Policy generalization across tasks
[…]
We observed the same effect when training with ActiveSigLIP on the large scale Webli dataset [6] (results summarised in Table 3). We trained reference models on both the raw 4B set as well as two extensively curated 1B / 350M subsets, which were then used to train subsequent learner models (Fig. 7). In all-but-one cases, filtering data with the 350M-trained referenced model produced the best results when transferring to 350M, 1B and 4B learner datasets. Notably, the significant gains observed over training with uniform sampling on the 350M subset suggest that our method can still improve over methods that pre-filter datasets once but then train with uniform samples [12].
4.7 Comparison to prior multimodal art
We leverage the insight that pre-training a reference model on clean data can facilitate learning on larger, noisy data for training our final model. Here, we train a reference model on LTIP, then use it to train a new model on the much larger mixture of LTIP and ALIGN, following [2], for a total of 3 and 8 billion training examples. Table 3 shows that in this training regime, ActiveCLIP surpasses models trained with significantly more data on ImageNet 0-Shot classification and COCO retrieval. Finally, we find that our active learning method is complementary to recent advances in multimodal learning: ActiveSigLIP significantly improves uniformly-trained SigLIP [44] in both COCO retrieval metrics.
@@ 4.7 Comparison to prior multimodal art
SigLIP 3B 72.1 60.7 42.7
ActiveSigLIP 3B 72.0 63.5 45.3
Table 3: Comparison of ActiveCLIP to public multimodal pre-taining methods. ActiveCLIP outperforms models trained with the same or more data (CLIP, [30]; EVA-CLIP, [37]; and OpenCLIP, [17]). ActiveSigLIP produced significant gains over the baseline SOTA performance of SigLIP [44]. benefits from additional ImageNet21K pretraining (+2B training examples). All ActiveCLIP/SigLIP models use a reference model trained on LTIP to guide learning on a mixture of ALIGN and LTIP.
@@ 5 Discussion
In this work, we have presented a new method for active data selection that builds upon and simplifies the concept of ‘learnability’. Our experiments demonstrate that this approach can significantly reduce the computation required for large-scale pretraining, compared to training with uniform samples. To our knowledge, this is the first active learning method that is more efficient than training with uniform samples when accounting for total FLOPs, and that does not rely on hand-designed features, allowing broad application across training setups. We have validated this by showing results on classification and contrastive pre-training, and found that our data selection policies continue to produce efficiency gains in the large-scale regime and can generalize effectively across task modalities. Collectively, our experiments also illustrate a Pareto frontier that allows trading off actor/data-selection computation against savings in training iterations, suggesting an alternative path to improved performance beyond scaling training batch sizes.
This work focused on supervised pre-training for images, but further work could involve extending our method to other modalities and training schemes such as language, video, and generative modeling. An important note is that all our experiments present results from filtering only 50% of the data; further gains may be possible by filtering more aggressively, at the cost of further overheads. In particular, aggressive data-selection coupled with efficient scoring schemes such as the ones proposed here could test the hypothesis that large-scale pretraining can benefit from exponential, rather than power-law, scaling behavior.
Appendix 0.A Supplementary
0.A.1 Details of total compute calculations